A 0.220-kg piece of aluminum that has a temperature of -191 °C is added to 1.5 kg of water that has a temperature of 2.4 °C. At equilibrium the temperature is 0 °C. Ignoring the container and assuming that the heat exchanged with the surroundings is negligible, determine the mass of water that has been frozen into ice.
Answer #1

The amount of heat added to the aluminum is m*c*deltaT = 0.22kg*900J/kgC*191deg C = 37818J

Therefore the water must give up 37818J.

This will be in two steps. The first step will be to cool all of the water to 0degC

This heat is m*c*deltaT = 1.5kg*4186J/kgC*2.4degC = 15069 J.

So in freezing the water must give up 37818-15069J = 22749 J

The latent heat of fusion for water is 3.34x10^5J/kg Use Q = m*Lf so m = Q/Lf = 22749J/3.34x10^5J/kg = 0.068 kg . So, 0.068 kg or 68 g of the water will freeze.